10 9 10 9 4.00 x 109 > m what is the electric field at location < 10 9 2.00 x 10-9 A π 4.00 pi minus particle, which has charge-e, s at location < 7.00 10.9 > m, due to the π-particle? 4.00 2.00 E= -Select- An antiproton (same mass as the proton, charge-e) is at the observation location. what is the force on the antiproton, due to the π Select--part 1 and 2^Lithium nucleus affected by an electric field A proton at location A makes an electric field E1 at location B. A different proton, placed at location B, experiences a force F Fi If E1800 N/C, what is F 1l? IF, 11 = Now the proton at B is removed and replaced by a lithium nucleus, containing three protons and four neutrons. The proton at location A remains in place What is the magnitude of the electric force on the lithium nucleus? F on Lil Now the Lithium nucleus is removed, and an electron is placed at location B. The proton at location A remains in place. What is the magnitude of the electric force on the electron? = one-l Which arrow below best indicates the direction of the force on the electron due to the electric field? -Select- j zero magnitudepart 3-6^
Answer #1

o vecton on is given by! 32 = 10-[(-2, 2,4) – (7,-4, -1)] = 1-4[-4, 6;8] 4 3.-=[tax-2) + (oxc)+ (8x8)] xt-axloei 9 1x1 KM3+82
Reversing of sign will be needed if the system hasn’t considered the negative sign for pi-minus particle.